OpenStudy (anonymous):

A woman drove part of a 185 mi trip at 50 mi/h and the rest at 55 mi/h. Find the distance she traveled at 50 mi/h if her total driving time was 3 h and 30 min.

OpenStudy (anonymous):

Let x hours be time she drove at 50mi/hr then (3.5-x) hours she drove at 55mi/hr....... Now, Total distance travelled = 185 mi

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

wait no how do i solve?

OpenStudy (anonymous):

Hope...this helps...))

OpenStudy (anonymous):

the only thing i am confused about is how you get 180-y as the distance

OpenStudy (anonymous):

we knw total distance is 180 so....if the first one travels y miles....then the next one would travel 180 - y

OpenStudy (anonymous):

Did u get it...

OpenStudy (anonymous):

isnt the total difference 185

OpenStudy (anonymous):

yes....i am sorry i didnt notice

OpenStudy (anonymous):

i hope..u understood wat i said

OpenStudy (anonymous):

wait so far i got y= 55x -7.5 is that correct?

OpenStudy (anonymous):

Let x hours be time she drove at 50mi/hr let the distance coverd During this time y S = D/T = 50 * x = y -------- 1 CaseII Time = 3.5 - x Distance = 180 - y Speed=55 S = D/T 55 (3.5 - x) = 185 - y 192.5 - 55x = 185- y -55x + y = -12.5 ---------2 nw u have two equations and two unknown variable.....find x and y

OpenStudy (anonymous):

i got -12.5 + 55x= 55x- 7.5

OpenStudy (anonymous):

Total distance: d = d1 + d2 = v1t + v2 ( 3.5 - t) = 55t + 50 ( 3.5 - t) @pepper0331 Do you understand until this part?

OpenStudy (anonymous):

185 mi = 5t + 175 mi

OpenStudy (anonymous):

Can you fill in the blanks now : -> t = (185 - 175) / 5 = .... Thus the distance d1: 55 * .... d2: 50 * ....

OpenStudy (anonymous):

Questions, comments ???

OpenStudy (anonymous):

yeah. i got the part about (185-175)/5 which is 2. but what is 2?? and how do you solve the answer with it?

OpenStudy (anonymous):

oh wait i just read the part above now let me solve

OpenStudy (anonymous):

wait is the answer 100. i did d=rt so d=100

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

Feel free to verify if the solution is unclear :)

OpenStudy (anonymous):

@pepper0331 the distance d2 Is NOT 100 mi Pay attention at: 55t + 50 ( 3.5 - t) With t = 2 hrs, the distance d1 at 55 mi/h : 55 * 2 = 110 mi the distance d2 at 50 mi/h : 50 ( 3.5 - 2 ) = 50 * 1.5 = 75 mi

OpenStudy (anonymous):

oh thanks